jacob has a certain amount of fencing he wants to use to make a dog run. if he makes the length 36 ft the width will be 9 ft, but if he makes the length 40 ft the width will be only 5 ft. write and equation in the form y= mx+b that can be used to find the width, y, for any given length x
Answer by josgarithmetic(39623)

So the perimeter is 90 ft.
2(36+9)=90
and
2(40+5)=90

x=length
y=width
2x+2y=90
divide both sides by 2
x+y=45
ANSWER:______y=-x+45________